OpenStudy (anonymous):

Find the zeros of the function by using the Quadratic Formula. I am terrible at math. f(x) = x2 + 3x - 7

hartnn (hartnn):

Compare your quadratic function with \(f(x)=ax^2+bx+c\) find a,b,c then the two zeros of the function are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

Thank you I got it.

hartnn (hartnn):

welcome ^_^